If a is 5 less than b use trial and improvement to find a and b so that a times b=2250
Answer by stanbon(75887)

If a is 5 less than b use trial and improvement to find a and b so that a times b=2250
-----
a = b-5
--------
Equation:
a*b = 2250
(b-5)b = 2250
b^2-5b-2250 = 0
-----
Factor::
(b+45)(b-50) = 0
----
b = -45 or b = 50
---
If a = -45, a = -50
If b = 50, a = 45
=========================
Cheers,
Stan H.
==================